graph the equation y+3=1/2(x+2)
Answer by mananth(16946)

y+3=1/2(x+2)
multiply equation by 2
2(y+3)=x+2
2y+6=x+2
2y=x+2-6
2y=x-4
/2
y=(1/2)x-2
y= 1/ 2 x + -2

Assume values of x and plug in the equation to get values of y.
Plot the points
x 0 2 6
y= -2 -1 1
(x,y) ( 0,-2),(2,-1),(6,1)